OpenStudy (maddiep5):

evaluate the dot product of (8,4) and (-3,5)

OpenStudy (dumbcow):

dot product ..... \[<u_1,u_2> * <v_1,v_2> = u_1 v_1 + u_2 v_2\]

OpenStudy (maddiep5):

never mind it is -4

OpenStudy (owlcoffee):

The dot product has a formula, and I believe what yo ustated here are two vectors, let's call them \(\vec u (8,4)\) and \(v(-3,5)\) respectively. The dot product of u and v is defined as: \[\vec u .\vec v =\left| \vec u \right|\left| \vec v \right| \cos \theta\] Where theta is the angle between them. Consecutively, we can also calculate the dot product as the sum of the interproduct of the coordinates, this of course derives from the formula above: \[\vec u . \vec v =(x_u)(x_v)+(y_u)(y_v)\]